### v3

# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

Abstract
Oversmoothing is a fundamental challenge in graph neural networks (GNNs): as the number of layers increases, node embeddings become increasingly similar, and model performance drops sharply. Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy. While these metrics are related to oversmoothing, we argue they have critical limitations and fail to reliably capture oversmoothing in realistic scenarios. For instance, they provide meaningful insights only for very deep networks and under somewhat strict conditions on the norm of network weights and feature representations. As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations. We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights. To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing in the nonlinear setting without assumptions on the boundedness of the weight matrices. Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures. Our results show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail. Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
1 Introduction
Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning [25, 39], personalised recommendation systems [7, 26], social network analysis [10], and 3D mesh classification [37]. Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a limit on their depth. This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing [19, 23].
Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants [24, 4, 2, 21, 8, 45, 30]. At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix. In linear GNNs without bias terms, this eigenspace is often known and easily computable via e.g. the power method. However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
Therefore, these metrics are often considered as sufficient but not necessary evidences for oversmoothing [33]. Despite this, there is a considerable body of literature using these unreliable metrics as their evidences for non-occurrence of oversmoothing in GNNs [47, 6, 35, 44, 20, 22, 34, 9, 36, 43, 29, 42].
As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed. Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable. GNNs [44, 35, 34, 45, 29, 42], where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero. Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
Furthermore, we note that these metrics can only indicate oversmoothing when their values converge exactly to zero, corresponding to either an exact alignment to the dominant eigenspace or to the feature representation matrix collapsing to the all-zero matrix. This double implication presents an issue: in realistic settings with a large but not excessively large number of layers, we may observe the decay of the oversmoothing metric by, say, two orders of magnitude while still being far from zero. In such cases, it is unclear whether the features are aligning with the dominant eigenspace, simply decreasing in magnitude, or exhibiting neither of the two behaviours. As a result, these types of metrics provide little to no explanation for the degradation of GNN performance.
As an alternative to address these shortcomings, we advocate for the use of a continuous approximation of the rank of the network’s feature representations to measure oversmoothing. Our experimental evaluation across various GNN architectures trained for node classification demonstrates that continuous rank relaxations, such as the numerical rank and the effective rank, correlate strongly with performance degradation in independently trained GNNs—even in settings where popular energy-like metrics show little to no correlation.
Overall, the main contributions of this paper are as follows:
  • We review popular oversmoothing metrics in the current literature and simplify their theoretical analysis from a novel perspective of nonlinear activation eigenvectors.
  • We propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
  • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights. To the best of our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
  • We provide extensive numerical evidence that continuous rank relaxation functions better capture oversmoothing than commonly used Dirichlet-like metrics.
2 Background
2.1 Graph Convolutional Network
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges. Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees: $A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2}$, where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix. The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph. At each layer $l$, the node feature update of Graph Convolutional Network (GCN) [17] follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
2.2 Graph Attention Network
Graph Attention Networks (GATs) [41, 3] perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$. The corresponding feature update is $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$
3 Oversmoothing
Oversmoothing can be broadly understood as an increase in similarity between node features as inputs are propagated through an increasing number of message-passing layers, leading to a noticeable decline in GNN performance. However, the precise definition of this phenomenon varies across different sources. Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other. This definition is motivated by the behaviour of a linear GCN: $X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}$. Indeed, if $\widetilde{A}$ is the adjacency matrix of a fully connected graph, $A$ will have spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will converge toward the eigenspace spanned by the dominant eigenvector. Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$ [40].
As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$. Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{\top}X^{(l)}\|$, converges to zero.
3.1 Existing Oversmoothing Metrics
Motivated by the discussion about the linear case, oversmoothing is thus typically quantified and analysed in terms of the convergence of some node similarity metrics towards zero. In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$. The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features [4, 33]
$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2},$ (1)
where $u_{i}$ is the $i$-th entry of the dominant eigenvector of the message-passing matrix. It thus immediately follows from our discussion on the linear setting that $E_{\mathrm{Dir}}(X^{(l)})$ converges to zero as $l\to\infty$ for a linear GCN with converging weights product $W^{(0)}\cdots W^{(l)}$. This intuition suggests that a similar behaviour may occur for “smooth-enough” nonlinearities.
In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and [4] have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$. In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero [45]
$E_{\mathrm{Proj}}(X)=\|X-\mathcal{P}X\|^{2}_{F}$ (2)
where $\mathcal{P}=uu^{\top}$ is the projection matrix on the space spanned by the dominant eigenvector $u$ of the matrices $A^{(l)}$. In particular, this metric can be used only if the dominant eigenvector $u$ is the same for all $l$; this is, for example, the case with row stochastic matrices or when $A^{(l)}=A$ for all $l$.
3.2 A Unifying Perspective Based on the Eigenvectors of Nonlinear Activations
In the following discussion, we present a unified and more general perspective of the necessary conditions to have oversmoothing in the sense of the classical metrics, based on the eigenvectors of a nonlinear activation function. In the interest of space, longer proofs for this and the subsequent sections are moved to Appendix A.
Definition 3.1.
We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}\in\mathbb{R}$ such that $\sigma(tu)=\mu_{t}u$.
With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metric $E_{\mathrm{Proj}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$. Specifically, Theorem 3.2 shows that this holds provided all matrices $A^{(l)}$ share a common dominant eigenvector $u$, which is also an eigenvector of $\sigma$.
This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature. For example, (a) the proof by [4] applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by [45] holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
Theorem 3.2.
Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A_{l}$ for any $l$ and is also an eigenvector of the activation function $\sigma$. If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, and $\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0$, then
$E_{\mathrm{Proj}}(X^{(L)})\to 0\quad\text{ as }\quad L\to\infty.$
Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$. Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by [4] as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$. Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$. Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$. If $\sigma=\otimes\psi$ is a nonlinear activation function acting entry-wise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$. Therefore, Theorem 3.2 implies that if the weights are sufficiently small, the features align independently of the activation function used. This is consistent with the results in [45]. However, we note that the bounds on the weights required by Theorem 3.2 and those in [45] on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant. Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs. In addition to offering a different and unifying theoretical perspective on the results in [4, 45], we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.
4 Energy-like Metrics: What Can Go Wrong
Energy-like metrics such as $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ are among the most commonly used oversmoothing metrics. However, they suffer from inherent limitations that hinder their practical usability and informational content.
One important limitation of these metrics is that they indicate oversmoothing only in the limit of infinitely many layers, when their values converge exactly to zero. Since they measure a form of absolute distance, a small but nonzero value does not provide any meaningful information. On the other hand, convergence to zero corresponds to either perfect alignment with the dominant eigenspace or the collapse of the feature representation matrix to the all-zero matrix. While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing. Moreover, this convergence property requires the weights to be bounded. However, in practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with arbitrarily large (or small) weight magnitudes. This aligns with our intuition and what occurs in the linear case. Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ is dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph. More precisely, the following theorem holds:
Theorem 4.1.
[…]
$\lim_{l\to\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0$
with a linear rate of convergence $|\lambda_{2}/\lambda_{1}|$.
In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely $X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{\top}+R(l)\Big{)}$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$. This results in weakly expressive feature representations, independently of the magnitude of the feature weights. This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, this is e.g. the case when $\lambda_{1}>1$ having $X^{(l)}\rightarrow\infty$.
Another important limitation of these metrics is their dependence on a specific, known dominant eigenspace, which must either be explicitly known or computed in advance. Consequently, their applicability is strongly tied to the specific architecture of the network. In particular, the dominant eigenvector $u$ of $A^{(l)}$ must be known and remain the same for all $l$. This requirement excludes their use in most cases where $A^{(l)}$ varies with $l$, such as when $A^{(l)}$ is the standard weighted adjacency matrix of the graph.
5 The Rank as a Measure of Oversmoothing
Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing [14]. However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure. A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank [31, 32, 1]. The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm. The numerical rank is given by $\text{NumRank}(X)=\|X\|_{F}^{2}/\|X\|_{2}^{2}$. Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$
where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps addressing the limitations of Dirichlet-like measures. Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one. A small value of the effective rank directly implies that the feature representations are low-rank, suggesting a potentially suboptimal network architecture.
In Figure 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned. This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one. In contrast, continuous rank functions provide a more reliable measure of approximate feature alignment. Later, in Figure 2, we demonstrate that the same phenomenon occurs in GNNs trained on real datasets, where exact feature alignment is rare. In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
# 1 # 2 # 3 # 4
Refer to caption Refer to caption Refer to caption Refer to caption
$E_{\mathrm{Dir}}$ 0 0 13.25 77.78
$E_{\mathrm{Proj}}$ 0 0 0.83 0.97
MAD 0 0.81 0.81 0.57
NumRank 1 1 1.01 1.78
Erank 1 1 1.36 1.99
Figure 1: Toy scenarios depicting the behaviour of oversmoothing metrics. Each plot contains 50 nodes, each with two features plotted on the x-y axis. The features are: # 1 of the same value; # 2 perfectly aligned with the same vector; # 3 aligned to the same vector except for one (red) point; # 4 sampled from a uniform distribution. MAD (Sec. 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 although features are oversmoothing by definition. $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing. To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, the first feature was considered in place of $u$ in (1) and (2).
5.1 Theoretical Analysis of Rank Decay
In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entry-wise nonnegative. While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
We begin with several useful observations. Let $u$ be the dominant eigenvector of $A$ corresponding to $\lambda_{1}$ and satisfying $\|u\|=1$. Consider the projection $\mathcal{P}=uu^{\top}$. Given a matrix $X$, we can decompose it as $X=\mathcal{P}X+(I-\mathcal{P})X$. Since $u$ is a unit vector, it follows that $\|\mathcal{P}\|_{2}=1$, and therefore,
$\|X\|_{2}=\|\mathcal{P}\|_{2}\|X\|_{2}\geq\|\mathcal{P}X\|_{2}.$ (3)
Moreover, since $\mathcal{P}X$ and $(I-\mathcal{P})X$ are orthogonal with respect to the Frobenius inner product, we have $\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}=\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}$. Thus, we obtain the following bound:
$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$ (4)
The above inequality, together with Theorem 4.1, allows us to establish the convergence of the numerical rank for linear networks.
The Linear Case
Consider a linear GCN of the form $X^{(l+1)}=AX^{(l)}W^{(l)}$, where $A$ has a simple dominant eigenvalue $\lambda_{1}$ satisfying $|\lambda_{1}|\geq|\lambda_{2}|$. We have already noted that $\|X\|_{2}\geq\|\mathcal{P}X\|_{2}$, meaning that the numerical rank converges to one if $\|(I-\mathcal{P})X\|_{F}/\|X\|_{2}$ decays to zero. This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction. As established in Theorem 4.1, this is almost surely the case in linear GNNs. As a direct consequence, we obtain the following result:
Theorem 5.1.
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Under the same assumptions as in Theorem 4.1, the following identity holds almost surely:
$\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial. In the next subsection, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
5.1.1 The Nonlinear Case
To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to [18, 12] and the reference therein for further details.
Assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$. On $\mathcal{K}$, it is possible to introduce the partial ordering
$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$
where $\bar{\mathcal{K}}$ denotes the closed cone of nonnegative vectors. Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$. Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$, $d_{H}(x,y)=\log(M(x/y)/m(x/y)).$ Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$. However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone. It is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
$d_{H}(Ax,Ay)\leq\beta d_{H}(x,y)\qquad\forall x,y\in\mathcal{K}$ (5)
with $0\leq\beta\leq 1$. The last inequality is a consequence of the fact that $A$ preserves the ordering induced by the cone, i.e. if $x\geq_{\mathcal{K}}y$, then $Ax\geq_{\mathcal{K}}Ay$. In particular, we recall that (5) always hold for $\beta<1$ when $A$ is entry-wise strictly positive, as positive matrices map the whole $\bar{\mathcal{K}}$ in its interior, i.e. $x\geq_{\mathcal{K}}y$ implies $Ax>Ay$. In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$. To this end, we recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
Definition 5.2.
A family of nonnegative irreducible matrices $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$
Note that the above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
Consider now first the case of a linear function that can be represented in the form $F(X)=AXW$, where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively. Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
Lemma 5.3.
Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$. Assume also $X$ to be strictly positive and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$. If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
$\max_{i}d_{H}(F(X)_{:,i},u)\leq\beta_{C}\max_{i}d_{H}(X_{:,i},u),$
where $Y_{:,i}$ denotes the $i$-th column of $Y$, and $\beta_{C}<1$.
Nonlinear Perron-Frobenius theory extends some of the results about nonnegative matrices to particular classes of nonlinear functions. A function $\sigma\in C(\mathcal{K},\mathcal{K})$ is order preserving if given any $x,y\in\mathcal{K}$ with $x\geq_{\mathcal{K}}y$ then $\sigma(x)\geq_{\mathcal{K}}\sigma(y)$. Moreover $\sigma$ is subhomogenenous if $\sigma(\lambda x)\leq_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and any $\lambda>1$. In particular, it is strictly subhomogeneous if $\sigma(\lambda x)\ll_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda>1$ and homogeneous if $\sigma(\lambda x)=\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda\geq 0$
Subhomogeneity is a useful property and of practical utility. In fact, as discussed in  [38, 27], a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$. For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
Theorem 5.4.
Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative, for any $l$. Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and that it is also an eigenvector of $\sigma$. Then, if the matrices $\{A^{(l)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$, it holds
$\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous. Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$. Then one can easily show that $\sigma$ is itself subhomogeneous on +$\mathbb{R}^{N}_{+}$. We have,
Proposition 5.5.
Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R}_{+},\mathbb{R}_{+})$ be order preserving. Then: 1) If $\sigma$ is homogeneous, any positive vector is an eigenvector of $\sigma$. 2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
As a consequence of the above result, we find that the numerical rank of the features collapses to one both for GCNs with LeakyReLU activation function (the setting of [4]) as well as for GATs with any kind of subhomogeneous activation function (the setting of [45]), as the constant vector is always an eigenvector of an entry-wise nonlinear map. For GCNs with any other subhomogeneous activation functions, the convergence of the numerical rank to 1 as stated in Theorem 5.1 may not hold as discussed in Appendix D.
Dataset Model $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
Standard Normalized Standard Normalized
Cora GCN -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2693
GAT -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2493
Citeseer GCN -0.8442 0.4350 -0.8913 -0.8667 -0.7169 0.9700 0.6795 0.4380
GAT -0.9576 0.0664 -0.9585 -0.9080 0.3722 0.9915 0.8047 0.4672
Pubmed GCN -0.9068 0.7006 -0.8508 -0.1109 0.6205 0.9464 0.9268 0.5225
GAT -0.8735 -0.3684 -0.8541 -0.4102 -0.3932 0.9270 0.9721 0.5564
Squirrel GCN -0.7774 0.4171 -0.7602 -0.3258 -0.8247 0.6316 0.9582 0.8457
GAT -0.6864 -0.5503 -0.7364 -0.7253 0.5002 0.8538 0.6840 0.7533
Chameleon GCN -0.9223 0.1504 -0.9163 -0.8201 -0.8809 0.9387 0.9014 0.6195
GAT -0.8721 0.1942 -0.9089 -0.8234 0.2803 0.9446 0.8799 0.6332
Amazon Ratings GCN -0.9297 0.8809 -0.9079 -0.3423 0.9201 0.9301 0.8049 0.8562
GAT -0.9388 0.5277 -0.9089 -0.1617 0.6545 0.9248 0.8764 0.8384
OGB-Arxiv GCN 0.7738 0.9194 0.5740 -0.2738 0.2822 0.9682 0.9091 0.0939
GAT -0.4097 0.9439 -0.7230 0.8985 0.8492 0.7740 0.9781 0.2310
Average correlation -0.7179 0.4036 -0.7571 -0.4504 0.1601 0.9103 0.8374
Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and of depths from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets. For Erank and NumRank, we subtract 1 so that both metrics approach zero. The rightmost column reports the ratio of classification accuracy between GNNs with 2 and 24 layers. Some heterophilic datasets may be more resilient to the increasing network depth, in-line with observations from the literature, e.g. [13]. Additional results on other datasets and activation functions are presented in Appendix E, and with additional network components in Appendix F.
6 Experiments
Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs [35, 44, 34, 45]. We emphasize that this is an overly simplified setting. In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe. The experiments that we present in this section validate the robustness of the relaxed rank metrics in quantifying oversmoothing in GNNs against the other metrics.
In particular, we compare how different oversmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data. In our experiments, we consider the following metrics:
  • The Dirichlet Energy $E_{\mathrm{Dir}}$ [4, 33] and its variant $E_{\mathrm{Proj}}$ [45]. Both are discussed in Section 3.1, see in particular (1) and (2).
  • Normalized versions of Dirichlet energy and its variant, $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$. Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features. Metrics with global normalization like the ones we consider here have also been proposed in [8, 30, 20].
  • The Mean Average Distance (MAD) [5]
    $\textstyle\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\Big{(}1-\frac{X_{i,:}^{\top}X_{j,:}}{|X_{i,:}||X_{j,:}|}\Big{)}.$
    It measures the cosine similarity between the neighbouring nodes. Unlike previous baselines, this oversmoothing metric does not take into account the dominant eigenvector of the matrices $A^{(l)}$.
  • Relaxed rank metrics: We consider the Numerical Rank and Effective Rank. Both are discussed in Section 5. We point out that from our theoretical investigation, in particular from (4), the numerical rank decays to $1$ faster than the decay of the normalized $E_{\mathrm{Proj}}$ energy to zero. This further supports the use of the Numerical Rank as an improved measure of oversmoothing with respect to $E_{\mathrm{Proj}}$.
Figure 2: Four examples of the metric behaviours computed at the last hidden layer of separately trained GCNs of increasing depths. For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$. In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$. Note that the effective rank and numerical rank of the input features $X^{(0)}$ are about 1084 and 13.6, respectively. Additional results are attached in Appendix G.
In Table 1 and Figure 2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits. We follow the standard setups of GCN and GAT as stated in Sections 2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function. For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained. The oversmoothing metric and accuracy results are averaged over 10 separately trained GNNs. All GNNs are trained with NAdam Optimizer and a constant learning rate of 0.01. The oversmoothing metrics are computed at the last hidden layer before the output layer. In Figure 2 and in Appendix G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth. These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics. By contrast, the rank of the feature representations decreases drastically, following closely the behaviour of the network’s accuracy. These findings are further sustained by the results shown in Table 1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model. The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
Extensions of these results are provided in Appendices E and F. In Appendix D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights. This study serves to validate our theoretical findings on the convergence of relaxed rank metrics and to demonstrate that such untrained settings offer little insight into the ability of existing metrics to quantify oversmoothing in realistic, trained networks.
7 Conclusion
In this paper, we have discussed the problem of quantifying oversmoothing in message-passing GNNs. After simplifying the existing theoretical analysis using nonlinear activation eigenvectors and discussing the limitations of the leading oversmoothing measures, we propose the use of the rank of the features as a better measure of oversmoothing. We provide extensive experiments to validate the robustness of the effective rank against the classical measures. In addition, we have proved theoretically the decay of the rank of the features for message-passing GNNs.
Acknowledgement
KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). PD was supported by the MUR-PRO3 grant STANDS. DJH was supported by a Fellowship from the Leverhulme Trust. FT was partially funded by the PRIN-MUR project MOLE, code 2022ZK5ME7, and PRIN-PNRR project FIN4GEO within the European Union’s Next Generation EU framework, Mission 4, Component 2, CUP P2022BNB97.
Appendix A Proofs of the main results
A.1 Proof of Theorem 3.2
We start proving that
$\|(I-\mathcal{P})X^{(l+1)}\|_{F}\leq\|(I-\mathcal{P})A^{(l)}X^{(l)}W^{(l)}\|_{F},$ (6)
where $\mathcal{P}=uu^{\top}/\|u\|^{2}$ is the projection matrix on the linear space spanned by $u$.
To this end, let $\pi:=\mathrm{span}\{uv^{\top}\,|\;v\in\mathbb{R}^{d}\}$ be the $1$-dimensional matrix subspace of the rank-$1$ matrices having columns aligned to $u$. Then it is easy to note that given some matrix $X$, $(I-\mathcal{P})X$ provides the projection of the matrix $X$ on the subspace $\pi$, i.e.
$(I-\mathcal{P})X=\mathrm{proj}_{\pi}(X).$ (7)
Indeed $\langle(I-\mathcal{P})X,uv^{\top}\rangle_{F}=\mathrm{Tr}(vu^{\top}(I-uu^{\top}/\|u\|^{2})X)=0$. In particular, since the projection realizes the minimal distance, we have that
$\|X-\mathcal{P}X\|_{F}\leq\|X-uv^{\top}\|_{F}\qquad\forall v\in\mathbb{R}^{d}.$ (8)
Now observe that $\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})=u\bar{v}^{\top}$ for some $\bar{v}$. Indeed, writing $v^{\top}=u^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)}$, we have that the $i$-th column of $\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})$ is equal to $\sigma(v_{i}u)=\bar{v}_{i}u$ for some $\bar{v}_{i}$, because $u$ is an eigenvector of $\sigma$. As a consequence we have
$\|(I-\mathcal{P})X^{(l)}\|_{F}$ $\leq\|X^{(l)}-\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$ (9)
$=\|\sigma(A^{(l-1)}X^{(l-1)}W^{(l-1)})-\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$
$\leq\|(I-\mathcal{P})A^{(l-1)}X^{(l-1)}W^{(l-1)}\|_{F}$
where we have used the $1$-Lipschitz property of $\sigma$. This concludes the proof of (6)
To conclude the proof of the theorem observe that, in the decomposition
$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}\mathcal{P}+(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
the matrix $(I-\mathcal{P})A^{(l)}\mathcal{P}$ is zero because $A^{(l)}u=\lambda_{1}^{l}u$ for any $l$. Thus
$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
and, from (6) and the inequality $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$, we have
$\|(I-\mathcal{P})X^{(L)}\|_{F}\leq\Big{(}\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}\Big{)}\|X^{(0)}\|_{F}.$
So the thesis follows from the hypothesis about the product $\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}$.
A.2 Proof for Theorem 4.1
Start by studying the norm of $(I-\mathcal{P})X^{(l)}$. Then looking at the shape of the powers of the Jordan blocks matrix it is not difficult to note that $\widetilde{T}^{l}=O({l\choose N}\lambda_{2}^{l-N})$ for $l$ larger than $N$. In particular if we look at the explicit expression of $(I-\mathcal{P})X^{(l)}$
$(I-P)X^{(l)}=\begin{pmatrix}0&(I-\mathcal{P})\tilde{M}\end{pmatrix}\begin{pmatrix}0&0\\ 0&\tilde{T}^{l}\end{pmatrix}M^{-1}X^{(0)}W^{(0)}\dots W^{(l-1)},$ (10)
we derive the upper bound
$\|(I-P)X^{(l)}\|_{F}\leq C{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F},$ (11)
for some positive constant $C$ that is independent on $l$.
Similarly we can observe that
$\|\mathcal{P}X^{(l)}\|_{F}\geq\|u^{\top}A^{l}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=$ (12)
$=\|\big{(}\lambda_{1}^{l}v_{1}^{\top}+u^{\top}\widetilde{M}O\Big{(}{l\choose N}\lambda_{2}^{l-N}\Big{)}\widetilde{M^{\prime}}\big{)}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\geq$
$\geq|\lambda_{1}|^{l}\Big{(}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big{\|}u^{\top}\widetilde{M}O\Big{(}{l\choose N}\Big{(}\frac{\lambda_{2}}{\lambda_{1}}\Big{)}^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big{\|}_{F}\Big{)}\geq$
$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}$
Now observe that under the randomness hypothesis from [11] and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system. In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
$c(\nu)-\epsilon\leq\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<c(\nu)+\epsilon$ (13)
i.e.
$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (14)
Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (15)
holding for any $l\geq l_{\epsilon}$ and any $w\in\{v_{1}\}\cup\{X_{0}^{\top}e_{i}\}_{i=1}^{N}$.
Next recall that $\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=\sqrt{\sum_{i}\|e_{i}^{\top}X^{(0)}W^{(0)}\|^{2}}$, meaning that almost surely, for $l\geq l_{\epsilon}$:
$Ne^{l(c(\nu)-\epsilon)}\leq\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\leq Ne^{l(c(\nu)+\epsilon)}.$ (16)
In particular for any $\epsilon$, there exists $l$ sufficiently large such that
($\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\right)$ (17)
and thus, since $|\lambda_{2}|<|\lambda_{1}|$ and we can choose $\epsilon$ arbitrarily small, almost surely it has limit equal to zero. In particular we can write
$\lim_{l}\frac{\|(I-\mathcal{P}X^{(l)})\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}\sim\lim_{l}\frac{{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}{|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}=0$ (18)
where we have used the same argument as before to state that the limit is zero.
A.3 An illustrative example
Consider the stochastic nonnegative primitive matrix
$A:=\begin{pmatrix}0&1\\[4.30554pt] 1/2&1/2\end{pmatrix}.$ (19)
Since $A$ is row stochastic, the dominant eigenvector is given by the constant vector $u=(1,1)$. In particular given a vector $x=(x_{1},x_{2})\in\mathcal{K}$ we have that
$d_{H}(x,u)=\log\Big{(}\frac{\max_{i=1,2}\{x_{i}\}}{\min_{i=1,2}\{x_{i}\}}\Big{)}=\begin{cases}\log\big{(}x_{1}/x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}x_{2}/x_{1}\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (20)
On the other hand $Ax=\big{(}x_{2}\,,\,(x_{1}+x_{2})/2\big{)}$, thus
$d_{H}(Ax,u)=\begin{cases}\log\big{(}(x_{1}+x_{2})/2x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}2x_{2}/(x_{1}+x_{2})\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (21)
Now observe that the set of vectors $x$ such that $d_{H}(x,u)\leq C$ is given by the points $\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}\cup\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$ with $C^{*}=e^{C}>1$. We want to prove the existence of some $\beta_{C}=1-\epsilon_{C}\in(0,1)$ such that $d_{H}(Ax,u)\leq\beta_{C}d_{H}(x,u)$ for any $x$ with $d_{H}(x,u)\leq C$.
To this end consider first the case of $x\in\Omega_{1}=\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}$. In this case
$d_{H}(Ax,u)=\log\Big{(}\frac{2x_{2}}{x_{1}+x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{1}$ (22)
$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}\leq\log\Big{(}\frac{x_{2}(x_{1}+x_{2})}{2x_{1}x_{2}}\Big{)}\qquad\forall x\in\Omega_{1}$
$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{1+t}{2}\Big{)}\qquad\forall t\in[1,C^{*}]$
$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{1+t}{2}\qquad\forall t\in[1,C^{*}]$
where we have written $t=x_{2}/x_{1}$. Second we consider the case of $x\in\Omega_{2}=\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$. In this case
$d_{H}(Ax,u)=\log\Big{(}\frac{x_{1}+x_{2}}{2x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{2}$ (23)
$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}\leq\log\Big{(}\frac{2x_{1}x_{2}}{x_{2}(x_{1}+x_{2})}\Big{)}\qquad\forall x\in\Omega_{2}$
$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{2t}{t+1}\Big{)}\qquad\forall t\in[1,C^{*}]$
$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{2t}{t+1}\qquad\forall t\in[1,C^{*}]$
where we have written $t=x_{1}/x_{2}$. Both the functions $t\rightarrow(1+t)/2$ and $t\rightarrow 2t/(t+1)$ are monotonically increasing for $t\in[1,\infty)$ with derivative $1/2$ in $t=1$, thus for any $C<\infty$ there exists some $10<\epsilon_{C}<$ 1 such that both (22) and (22) are satisfied in the interval $t\in[1,C^{*}]$. This shows that $A$ is contractive with respect to $u$ even if $A$ is not strictly positive.
A.4 proof of Lemma 5.3
By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{:,i},u\big{)}\leq C$, then
$d_{H}\Big{(}(AX)_{:,i},u\Big{)}=d_{H}\Big{(}(AX)_{:,i},\lambda_{1}(A)u\Big{)}=d_{H}\Big{(}(AX)_{:,i},Au\Big{)}\leq\beta d_{H}\big{(}X_{:,i},u\big{)}\qquad\forall i.$ (24)
for some $\beta<1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance.
Then note that, for any $i$, we can write $F(X)_{:,j}$ as follows
$\big{(}F(X)\big{)}_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (25)
Thus we CLAIM that given $x_{1},x_{2},y\in\mathcal{K}$ then
$d_{H}(x_{1}+x_{2},y)\leq\max\{d_{H}(x_{1},y),d_{H}(x_{2},y)\}.$ (26)
Observe that if the claim holds we have conlcuded the proof, indeed by induction it can trivially be extended from $2$ to $d$ points yielding
$d_{H}\Big{(}F(X)_{:,j},u\Big{)}\leq\max_{i}d_{H}\Big{(}W_{ij}(AX)_{:,i},u\Big{)}\leq\max_{j}d_{H}\Big{(}(AX)_{:,j},u\Big{)}\leq\beta\max_{j}d_{H}\Big{(}X_{:,j},u\Big{)},$ (27)
where we have used the scale-invariance property of the Hilbert distance and the fact that $\max_{i}W_{ij}>0$ for all $j$.
We only miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
$d_{H}(x_{1}+x_{2},y)$ $=\log\bigg{(}\sup_{j}\sup_{i}\frac{(x_{1})_{i}+(x_{2})_{i}}{(y)_{i}}\frac{(y)_{j}}{(x_{1})_{j}+(x_{2})_{j}}\bigg{)}$ (28)
$\leq\log\bigg{(}\sup_{i}\sup_{j}\max_{x_{1},x_{2}}\bigg{\{}\frac{(x_{1})_{i}}{(x_{1})_{j}},\frac{(x_{2})_{i}}{(x_{2})_{j}}\bigg{\}}\frac{(y)_{j}}{(y)_{i}}\bigg{)}$
$=\max_{x_{1},x_{2}}\big{\{}d_{H}(x_{1},y),d_{H}(x_{2},y)\big{\}}.$
concluding the proof.
A.5 proof of Theorem 5.4
To prove the theorem we start proving that, a continuous subhomogeneous and order-preserving function $\sigma$ with an eigenvector $u$ in the cone, is not nonexpansive in Hilbert distance with respect to $u$. Formally we claim that
$d_{H}\big{(}\sigma(y),u\big{)}\leq d_{H}\big{(}y,u\big{)}\qquad\forall y\in\mathcal{K}.$ (29)
To prove it let $y\in\mathcal{K}$ and assume w.l.o.g. that $\|y\|_{1}=t>0$ and $\|u\|_{1}=1$, then
$M(y/tu)=\max_{i=1,\dots,N}\frac{y_{i}}{t(u)_{i})}\geq\frac{\|y\|_{1}}{t\|u\|}=1\qquad m(y/tu)=\min_{i=1,\dots,N}\frac{y_{i}}{t(u)_{i}}\leq\frac{\|y\|_{1}}{t\|u\|_{1}}=1.$ (30)
By definition given $x,y\in\mathcal{K}$, $m(y/x)x\leq_{\mathcal{K}}y\leq_{\mathcal{K}}M(y/x)x$. Moreover we recall that since $u$ is an eigenvector for any $t>0$ there exists $\lambda_{t}>0$ such that $\sigma(tu)=\lambda_{t}u$. Thus we use the subhomogeneity of $\sigma$ and the fact that $u$ is an eigenvector of $\sigma$ to get the following inequalities:
$m(y/tu)\lambda_{t}tu\leq_{\mathcal{K}}\sigma\big{(}m(y/tu)tu\big{)}\leq_{\mathcal{K}}f(y)\leq_{\mathcal{K}}f\big{(}M(y/tu)tu\big{)}\leq_{\mathcal{K}}M(y/x_{c})\lambda_{t}tu.$ (31)
In particular we have $m(f(y)/tu)\geq\lambda_{t}m(y/tu)$ and $M(f(y)/tu)\leq\lambda_{t}M(y/tu)$. Finally the last inequalities and the scale invariance property of $d_{H}$ yield the thesis:
$d_{H}(f(y),u)=d_{H}(f(y),tu)=\log\Big{(}\frac{M(f(y)/tu)}{m(f(y)/tu)}\Big{)}\leq\log\Big{(}\frac{M(y/tu)}{m(y/tu)}\Big{)}=d_{H}(y,tu)=d_{H}(y,u)\,.$ (32)
Now using the claim above and Lemma 5.3, we can easily conclude that
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (33)
Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{:,i},u)$, then there exists some $\beta_{C}<1$ such that
$\max_{i}d_{H}(X^{(l)}_{:,i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{:,i},u).$ (34)
Finally to conclude we prove that, as a consequence of (33),
$\lim_{l\rightarrow\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0.$ (35)
As a consequence of the inequality:
$1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}},$ (36)
we have that (35) finally yields the thesis. To prove (35), we recall from Lemma 2.5.1 in [18] that for any $w$ such that $u^{\top}w=c$
$\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{T}(w,\mathcal{P}w)}-1),$ (37)
where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e. $\|x\|_{u}=u^{\top}x$ for any $x$ in the cone. In practice the norm induce by $u$ $\|\cdot\|_{u}$ can be defines by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{\top}x\leq 1\}$
Then since $\|\mathcal{P}w\|_{u}=\|w\|_{u}=u^{\top}w$, we have that $M(w/\mathcal{P}w)\geq 1$ and $m(w/\mathcal{P}w)\leq 1$. Thus $d_{T}(w,\mathcal{P}w)\leq d_{H}(w,\mathcal{P}w)$, yielding
$\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{H}(w,\mathcal{P}w)}-1).$ (38)
From the equivalence of the norms there exists some constant $c>0$ such that we can equivalently write
$\|w-\mathcal{P}w\|_{2}\leq C\|\mathcal{P}w\|_{2}(e^{d_{H}(w,\mathcal{P}w)}-1).$ (39)
Now recall that the squared frobenius norm of a matrix is the sum of the squared 2-norms of the its columns, so we can apply the last inequality to the matrix $X^{(l)}$ columnwise obtaining:
$\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})\}}-1\big{)}^{2}.$ (40)
the proof is concluded using (33) and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding:
$1\leq\big{(}\mathrm{NumRank}(X^{(l)})\big{)}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},u)\}}-1\big{)}^{2}$ (41)
and concluding the proof.
A.6 proof of Proposition 5.5
We start from the homogenous case. Note that since $\psi$ is homogeneous we have that necessarily $f(t)=ct$ for all $t,c\geq 0$, this in particular means that every $u\in\mathbb{R}^{N}_{+}$ is an eigenvector of $\sigma$ with corresponding eigenvalue $\lambda_{1}=c$.
Then we can consider the subhomogeneous case. Assume that we have $u\in\mathbb{R}_{+}^{N}$ that is an eigenvector of $\sigma$ with eigenvalue $\lambda$ and $u_{i}>0$ for all $i$, then
$\psi(u_{i})=\lambda u_{i}\qquad\forall i=1,\dots,N.$ (42)
By strict subhomogeneity this means that necessarily $u$ is constant, indeed if $u_{i}>u_{j}>0$ then
$\lambda u_{j}=\psi(u_{j})=\psi\Big{(}u_{j}\frac{u_{i}}{u_{i}}\Big{)}>\frac{u_{j}}{u_{i}}\psi(u_{i})=\lambda u_{j},$ (43)
yielding a contradiction. In particular any constant vector $u$ in $\mathbb{R}_{+}^{N}$ is easily proved to be an eigenvector of $\sigma$ relative to the eigenvalue $\lambda=\|\sigma(u)\|_{1}/\|u\|_{1}=\psi(u_{i})/u_{i}$ where $u_{i}$ is any entry of $u$.
Appendix B Limitations
Although a small effective rank or numerical rank indicates oversmoothing and can be subsequently linked to the underperformance of GNNs, a large effective rank or numerical rank does not necessarily correspond to a good network performance. Prior study has suggested some degree of smoothing can be beneficial [16], and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern. We note that this limitation is not specific to the two relaxed rank measures, as the same argument is directly applicable to all other oversmoothing metrics.
Consequently, as shown in Table 4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs. This in-turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
Appendix C Computational Complexity Analysis
Let $N$, $D$, $E$ the number of nodes, features and edges of a graph $\mathcal{G}$, the computational cost of the Dirichlet energy is $O(E\times D)$, while the cost of the Projection Energy is $O(N\times D)$. In contrast, most of the computational cost for the numerical rank and the effective rank is given by the computation of the spectral radius of $X$, and of all the singular values of $X$ for the effective rank, respectively.
Standard results show that it is always possible to compute the full singular value decomposition (SVD), and subsequently the spectral radius, in $O(N\times D\times\min\{N,D\})$. However, in typical cases, the latter cost can be drastically reduced using different strategies. Firstly, in the case of numerical rank, the computational cost of the 2-norm is generally much smaller than the cost of the full SVD. Indeed, using Lanczos or power methods to compute it, the cost scales as $O(N\times D)$, and the methods converge typically very fast. In addition, both the effective rank and the numerical rank can be efficiently controlled by computing only the $k$-largest singular values of $X$. In particular, a truncated SVD containing the $k$-largest singular values can be computed in $O(N\times D\times k)$ using either deterministic algorithms or randomized SVD methods.
In general, the computational cost of metrics to quantify oversmoothing is marginal compared to the cost of training. In production, measuring the emergence of oversmoothing is done by training the model and checking the performance on the validation and test sets as the number of layers changes. The cost of additionally computing effective or numerical ranks is marginal. Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than $10~{}ms$ for a graph (or a subsection of it) with less than 2000 nodes.
Appendix D Additional Experimental Results with Synthetic Weights
In this section, we conduct an asymptotic ablation study using randomly sampled (synthetic) untrained weights. The aim of these experiments is twofold:
  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature [44, 35, 34, 45, 29, 42], as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems 5.1 and 5.4.
We construct a 10-node Barabasi-Albert graph with each node having 32 features. The weights are either an identity matrix or randomly sampled at each layer from a uniform distribution $\mathcal{U}(0,s)$, where $s$ depends on the settings: small weights ($s=0.05$) lead to an exponentially decaying $\|X^{(l)}\|_{F}$, and large weights ($s=0.1$) lead to an exploding $\|X^{(l)}\|_{F}$ for uncapped activation functions. For identity weights, $\|X^{(l)}\|_{F}$ is roughly constant (LReLU) or slowly decaying (Tanh). The feature initialization $X^{(0)}$ is sampled from $\mathcal{U}(0,1)$, and is iterated over 300 layers.
In this asymptotic synthetic setting, as presented in Table 2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse. However, this behaviour stands in stark contrast to results on trained networks presented in Tables 1 and 3 and Appendix G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
Moreover, these asymptotic results validate Theorems 5.1 and 5.4, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions. Without making any additional assumption on the normalization of the adjacency matrix, the effective rank and numerical rank do not generally decay to one when subhomogeneous activation functions, e.g. Tanh, are used in GCNs.
Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank
standard normalized standard normalized
GCN+LReLU+identity weights
GCN+Tanh+identity weights
GAT+LReLU+identity weights
GAT+Tanh+identity weights
GCN+LReLU+small weights
GCN+Tanh+small weights
GAT+LReLU+small weights
GAT+Tanh+small weights
GCN+LReLU+large weights
GCN+Tanh+large weights
GAT+LReLU+large weights
GAT+Tanh+large weights
Table 2: Additional results on very deep (300 layers) synthetic networks with randomly sampled weights. For Erank and NumRank, we subtract 1 so that both metrics converge to zero. ✓ indicates a decay of the corresponding metric to zero, ✗ indicates otherwise. Note that GAT has similar asymptotic behaviour to GCN with adjacency normalization $D^{-1}\widetilde{A}$.
Appendix E Additional Experimental Results on Activation Functions and Different Datasets
We extend Table 1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section 6.
Dataset Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
Standard Normalized Standard Normalized
Cora GCN+LReLU -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2693
GCN+Tanh 0.5243 0.9403 0.8610 0.9768 0.9734 0.9923 0.9784 0.1937
GAT+LReLU -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2493
GAT+Tanh 0.8300 0.8501 0.8676 0.9066 0.8603 0.9600 0.9515 0.1900
Citeseer GCN+LReLU -0.8442 0.4350 -0.8913 -0.8667 -0.7169 0.9700 0.6795 0.4380
GCN+Tanh 0.3420 0.8957 0.4631 0.9045 0.9605 0.9906 0.9457 0.3509
GAT+LReLU -0.9576 0.0664 -0.9585 -0.9080 0.3722 0.9915 0.8047 0.4672
GAT+Tanh 0.9234 0.8949 0.8276 0.8997 0.9176 0.9287 0.9024 0.3045
Pubmed GCN+LReLU -0.9068 0.7006 -0.8508 -0.1109 0.6205 0.9464 0.9268 0.5225
GCN+Tanh -0.2330 0.2657 0.2029 0.9137 0.8745 0.9328 0.9940 0.3883
GAT+LReLU -0.8735 -0.3684 -0.8541 -0.4102 -0.3932 0.9270 0.9721 0.5564
GAT+Tanh 0.2977 0.8411 0.7160 0.9331 0.8546 0.9303 0.8551 0.4464
Squirrel GCN+LReLU -0.7774 0.4171 -0.7602 -0.3258 -0.8247 0.6316 0.9582 0.8457
GCN+Tanh 0.6026 0.7736 0.1689 0.9377 0.7727 0.9680 0.9837 0.8152
GAT+LReLU -0.6864 -0.5503 -0.7364 -0.7253 0.5002 0.8538 0.6840 0.7533
GAT+Tanh -0.3606 -0.6557 -0.0363 0.8714 -0.7033 0.8911 0.8861 0.9103
Chameleon GCN+LReLU -0.9223 0.1504 -0.9163 -0.8201 -0.8809 0.9387 0.9014 0.6195
GCN+Tanh 0.2742 0.8796 -0.2541 0.8492 0.9272 0.9869 0.9841 0.7093
GAT+LReLU -0.8721 0.1942 -0.9089 -0.8234 0.2803 0.9446 0.8799 0.6332
GAT+Tanh 0.4090 0.5699 0.0743 0.8230 0.6006 0.9613 0.9143 0.7052
Amazon Ratings GCN+LReLU -0.9297 0.8809 -0.9079 -0.3423 0.9201 0.9301 0.8049 0.8562
GCN+Tanh -0.6960 -0.6289 -0.6910 0.8576 0.9423 0.9871 0.9327 0.8574
GAT+LReLU -0.9388 0.5277 -0.9089 -0.1617 0.6545 0.9248 0.8764 0.8384
GAT+Tanh 0.8354 0.8507 -0.6595 0.9245 0.9418 0.8954 0.8883 0.8382
Roman Empire GCN+LReLU -0.5635 0.7703 -0.6772 0.2575 0.6420 0.5833 0.5368 0.3891
GCN+Tanh 0.8018 0.9225 -0.6808 0.8359 0.8124 0.8570 0.8090 0.4067
GAT+LReLU -0.9174 0.5390 -0.9407 0.1868 0.7582 0.7221 0.8722 0.3705
GAT+Tanh 0.5767 0.5844 -0.4716 0.8819 0.7263 0.7332 0.8589 0.3652
OGB-Arxiv GCN+LReLU 0.7738 0.9194 0.5740 -0.2738 0.2822 0.9682 0.9091 0.0957
GCN+Tanh -0.6409 0.7041 -0.6808 0.9494 0.9487 0.9900 0.9876 0.1204
GAT+LReLU -0.4097 0.9439 -0.7230 0.8985 0.8492 0.7740 0.9781 0.2310
GAT+Tanh 0.7834 0.7896 -0.9277 0.9393 0.8222 0.7671 0.7861 0.2376
Average correlation -0.1956 0.5137 -0.3886 0.2669 0.4960 0.9007 0.8684
Table 3: Additional correlation coefficient results on homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings, Roman Empire) and large-scale (OGB-Arxiv) dataset.
Appendix F Additional Experimental Results with Different Network Components
Prior literature indicates that when adding additional components, such as bias [33] or residual terms [36], the Dirichlet energy does not decay. Therefore, we compare the correlation coefficients in Table 4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm [46], DropEdge [28, 15] and residual connections [36].
All experiments follow the setup described in Section 6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}$
Table 4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics. This finding confirms their superior consistency in detecting oversmoothing across a variety of architectural variants.
Furthermore, we note that when oversmoothing is effectively alleviated, e.g. when residual connection is used, all metrics have a poor correlation with the classification accuracy. This generic limitation is discussed in Appendix B.
Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
Standard Normalized Standard Normalized
GCN+LReLU+Bias -0.9300 0.7505 -0.9414 -0.3642 -0.2828 0.9847 0.7833 0.2110
GCN+Tanh+Bias 0.4086 0.8950 0.7479 0.9026 0.8996 0.9926 0.9814 0.2133
GCN+LReLU+LayerNorm -0.9132 0.8445 -0.9424 0.5591 0.9753 0.9736 0.9736 0.4882
GCN+Tanh+LayerNorm -0.7119 0.2560 -0.2069 0.9716 0.8695 0.9576 0.9684 0.1886
GCN+LReLU+BatchNorm 0.8789 0.7300 0.8872 0.7175 0.5094 0.6005 0.6577 0.8761
GCN+Tanh+BatchNorm -0.6033 0.2883 -0.6587 0.7717 0.4984 0.7377 0.7038 0.8157
GCN+LReLU+PairNorm -0.8165 0.3731 -0.8106 0.3885 0.4850 0.5597 0.6244 0.8556
GCN+Tanh+PairNorm -0.5963 0.0346 -0.5401 -0.3358 -0.0631 0.9838 0.9298 0.3123
GCN+LReLU+DropEdge -0.7497 0.7185 -0.7939 -0.7974 -0.4619 0.9720 0.5782 0.2515
GCN+Tanh+DropEdge 0.2319 0.8388 0.5872 0.8887 0.8763 0.9934 0.9558 0.2096
GCN+LReLU+Residual -0.0857 -0.3072 -0.0559 -0.1611 -0.1656 -0.2466 -0.1296 1.0046
GCN+Tanh+Residual -0.6751 -0.7406 -0.3031 -0.4019 -0.6043 -0.5425 -0.4897 1.0122
Average correlation -0.3801 0.3901 -0.2525 0.2616 0.2946 0.6638 0.6280
Table 4: Additional correlation coefficient results on GCNs with different network components.
Appendix G Metric Behaviour Examples on Cora, Citeseer and Pubmed
We extend Figure 2 to additional datasets and network components. The experimental setting is consistent with that of Section 6.
GCN + LReLU $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$
         [Uncaptioned image]          [Uncaptioned image]
GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.454981,\quad r^{*}_{\mathrm{NR}}=1.020319$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.858602,\quad r^{*}_{\mathrm{NR}}=1.021634$
         [Uncaptioned image]          [Uncaptioned image]
GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.017424,\quad r^{*}_{\mathrm{NR}}=1.000101$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.528688,\quad r^{*}_{\mathrm{NR}}=1.003682$
         [Uncaptioned image]          [Uncaptioned image]
GAT + LReLU $r^{*}_{\mathrm{ER}}=1.295623,\quad r^{*}_{\mathrm{NR}}=1.079250$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.002191,\quad r^{*}_{\mathrm{NR}}=1.000006$
         [Uncaptioned image]          [Uncaptioned image]
Table 5: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Cora Dataset. This table is an extension of figure 2
GCN + LReLU $r^{*}_{\mathrm{ER}}=1.750563,\quad r^{*}_{\mathrm{NR}}=1.142120$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.436012,\quad r^{*}_{\mathrm{NR}}=1.003035$
         [Uncaptioned image]          [Uncaptioned image]
GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.526297,\quad r^{*}_{\mathrm{NR}}=1.045451$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.752850,\quad r^{*}_{\mathrm{NR}}=1.007417$
         [Uncaptioned image]          [Uncaptioned image]
GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.006405,\quad r^{*}_{\mathrm{NR}}=1.000009$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.473417,\quad r^{*}_{\mathrm{NR}}=1.002555$
         [Uncaptioned image]          [Uncaptioned image]
GAT + LReLU $r^{*}_{\mathrm{ER}}=1.541711,\quad r^{*}_{\mathrm{NR}}=1.148837$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.003797,\quad r^{*}_{\mathrm{NR}}=1.000005$
         [Uncaptioned image]          [Uncaptioned image]
Table 6: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Citeseer Dataset.
GCN + LReLU $r^{*}_{\mathrm{ER}}=1.420913,\quad r^{*}_{\mathrm{NR}}=1.085279$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.838982,\quad r^{*}_{\mathrm{NR}}=1.014851$
         [Uncaptioned image]          [Uncaptioned image]
GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.443576,\quad r^{*}_{\mathrm{NR}}=1.018361$
         [Uncaptioned image]          [Uncaptioned image]
GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.002088,\quad r^{*}_{\mathrm{NR}}=1.000074$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.668289,\quad r^{*}_{\mathrm{NR}}=1.009275$
         [Uncaptioned image]          [Uncaptioned image]
GAT + LReLU $r^{*}_{\mathrm{ER}}=1.577163\quad r^{*}_{\mathrm{NR}}=1.119213$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.016096,\quad r^{*}_{\mathrm{NR}}=1.000019$
         [Uncaptioned image]          [Uncaptioned image]
Table 7: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Pubmed Dataset.